OpenStudy (sammyalabamy):

Evaluate the expression: v ⋅ w Given the vectors: r = <8, 8, -6>; v = <3, -8, -3>; w = <-4, -2, -6>

OpenStudy (nincompoop):

do you know how to dot product?

OpenStudy (abdullahm):

Here is how you find the dot product: \(x = <a_1, b_1, c_1>\) \(y = <a_2, b_2, c_2>\) \(\ x \cdot y = (a_1 \times a_2) + (b_1 \times b_2)+(c_1 \times c_2)\)

OpenStudy (sammyalabamy):

whats the third one for?

OpenStudy (abdullahm):

Which third one?

OpenStudy (sammyalabamy):

the r

OpenStudy (abdullahm):

They gave you the r most likely because there are other questions that will have to do with r

OpenStudy (sammyalabamy):

This is the final question and i am still lost. I did the (a1*a2)+(b1*b2)+(c1*c2) but what do i do after?

OpenStudy (nincompoop):

perhaps, read this so you're not just mechanically performing dot product without knowing what you're doing http://physics.info/vector-multiplication/

OpenStudy (sammyalabamy):

so do i do r+(v*w) or r*(v*w)

OpenStudy (abdullahm):

They only want the dot product of v and w

OpenStudy (nincompoop):

have you learnt about matrices before?

OpenStudy (sammyalabamy):

no i dont think so

OpenStudy (kainui):

The r vector is a red herring. Don't let it distract you, v ⋅ w does in no way depend on r at all since v and w are completely independent of r.